OpenStudy (anonymous):

I need help, simplistically, step by step how to solve 4x+6y-8z=1

OpenStudy (anonymous):

you must be missing some imformation... solve for a variable or what? we cannot solve with just that information given

OpenStudy (anonymous):

there is no step to solve it

OpenStudy (amistre64):

3 variables needs 3 equations..or is it 4 equations? its 3

OpenStudy (anonymous):

maybe all 3 together? 1) 4x+6y-8z=1 2) -6x-9y+12z=0 3) x-2y-2z=3

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

I still need help as simple as possible to solve it, I'm not good at math and don't remem ber how please

OpenStudy (anonymous):

first thing you need to do, is use the three equations to solve for a certain value of eachother, this is done by solving for a variable, then substituting

OpenStudy (anonymous):

i would start with number 3, we can say that x-2y-2z=3 can be rewritten as x=3+2y+2z (moved em all over to make x alone

OpenStudy (anonymous):

put it in a matrix calculator =p

OpenStudy (anonymous):

so now you have what x is, you can then put that in for the other equations: 1)4(3+2y+2z)+6y-8z=1 2) -6(3+2y+2z)-9y+12z=0

OpenStudy (anonymous):

... matricies would be easier, if you know how to do them?

OpenStudy (anonymous):

I don't know, I haven't attempted to do this kind of math in over 15 years

OpenStudy (anonymous):

the concept of matrices is a bit intuitive, if you want a shortcut to just the answer, a calculator would do it, if you need to know how to get it then ill continue algebraically

OpenStudy (anonymous):

how do you know what x is when y and z are still unknowns?

OpenStudy (anonymous):

please, if you don't mind

OpenStudy (anonymous):

we dont technically know x, but we know how to phrase x in terms of y and z, this way we can put that into the other equations as i did here: 1) 4(3+2y+2z)+6y-8z=1 2) -6(3+2y+2z)-9y+12z=0

OpenStudy (anonymous):

see how i first figured out (with equation 3) that x = 3+2y+2z... then put that new equation in for x in the other two equations? this way we no longer have three variables, we have two!

OpenStudy (anonymous):

Oh okay I understand how you came up with 1 and 2

OpenStudy (anonymous):

the goal is to solve for only one variable at a time, then find the other two after that

OpenStudy (anonymous):

x =3+2y+2z 4(3+2y+2z) +6y-8z =1 --> 12+8y+8z+6y-8z=1 --> 14y= -11 --> y = -11/14 x =3+2y+2z --> x= 3+2(-11/14)+2z --> x= (3-11/7)+2z x= (10/7)+2z -6x-9y+12z=0---> -6(10/7)+2z-9(-11/14)+12z=0 ----> -60/7 + 99/14 +14z=0 -----> (-120+99)/14 +14z=0 -------> z = (21/14)/14 z= ........ u got y and z, and plug in to x =3+2y+2z, u get x

OpenStudy (anonymous):

^there ya go haha

OpenStudy (anonymous):

You are awesome Ty

OpenStudy (anonymous):

ummm...I need help again, how does 12+8y+8z+6y-8z=1 become 14y=-11? I understand that when you multiple a positive with a negative it become positive, and I understand that 8z-8z cancels each other out, but then wouldn't it become 12+14y=1?

OpenStudy (anonymous):

OOOOOH nm, I get it, 12+14y=1 becomes 14y=1-12

OpenStudy (anonymous):

how does 3+2(-11/14) become 3-11/7?

OpenStudy (anonymous):

14:2 = 7 ---> 3 + - 11/7 ---> 3-11/7

OpenStudy (anonymous):

I'm sorry, I still don't understand how you come up with that. why does 2 do something with 14 and not -11?

OpenStudy (anonymous):

2(-11/14)= -22/14 = 11/7, if u want it stay -22/14 that is ok